dingas2002:

***PLEASE HELP PLS *** *Round to 1 decimal place 1. What is the mass (in kg) of a 2714 N ​offensive lineman? 2. What is the magnitude of the weight (in Newtons) of a 34.8 kg dog? 3. A 366 g water bottle sits on top of your desk. What is the magnitude of the Force of Gravity (in Newtons) that the water bottle exerts on the Earth?​ 4. A grandpa is doing pull ups with a tiger tied around his ankles. What is the total number of forces acting on the grandpa if he is hanging at rest from the pull up bar and the tiger is not touching the ground.

justjm:

1. Apply F=ma/ W=mg W = 2714 N m = in kg you must find g = 9.81 m/s^2 2. Apply W=mg W in N = 34.8 kg * 9.81 m/s^2\ 3. Apply W=mg again. Remember to convert 366 g to kg W = you must find in N m = 366 g, convert to kg!!! g = 9.81 m/s^2 4. Pretty rough for a grandpa, no? Maybe he should do something less risky...? Make a systems schema to help you understand the forces between bodies. Then make an FBD (force distribution diagram) and do a summation of forces in like directions. I'll list the forces. -Weight of grandpa -Weight of Tiger -Tension of Tiger on grandpa -Tension of the pull up bar including tiger and grandpa -Normal force of the pull up bar including grandpa and tiger -Normal force on grandpa from tiger